OpenStudy (anonymous):

Multiply both sides of the following equation by r, and use the fact that r^2=x^2 + y^2 to rewrite the equation in terms of x and y. r=-6sinø

jimthompson5910 (jim_thompson5910):

multiply both sides by r keep in mind that y = r*sin(theta)

jimthompson5910 (jim_thompson5910):

r=-6sin(theta) r*r=r*(-6sin(theta)) r^2=-6*(rsin(theta)) I'll let you finish

OpenStudy (anonymous):

So r^2 = -6y r= √-6y?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

use r^2 = x^2 + y^2 and r*sin(theta) = y

OpenStudy (anonymous):

Ah. -6y=x^2 + y^2 y=x^2 + y^2/-6

jimthompson5910 (jim_thompson5910):

-6y=x^2 + y^2 is correct try to convert it into the form (x-h)^2 + (y-k)^2 = r^2